Circular No. 8332

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 H3 (LARSEN) J. A. Larsen, Lunar and Planetary Laboratory, University of Arizona,

reports the discovery of a comet on Spacewatch CCD images taken with the 0.9-m f/3 reflector at Kitt Peak (discovery observation below), reporting a 17′′ tail in p.a. 280^{o}. Following posting on the NEO Confirmation Page, J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector) reported that his stacked 3-min CCD frames taken on Apr. 24.22 UT show a soft coma with diameter 6′′ and a 8′′ tail in p.a. 270^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Apr. 22.32787  14^{h}48^{m}21.^{s}99  −17^{o}10′23′′.4  19.2

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2004-H66.

T = 2004 Feb. 29.472 TT  ω = 343.^{o}313
e = 0.39907  Ω = 221.067 }2000.0
q = 2.50394 AU  i = 26.845
a = 4.16676 AU n^{o} = 0.115880 P = 8.51 years

SUPERNOVA 2004am IN M82 R. J. Beswick, T. W. B. Muxlow, M. K. Argo, and A. Pedlar, Jodrell

Bank Observatory, report multi-frequency upper limits on the early radio emission from SN 2004am (cf. IAUC 8297, 8299), as part of their ongoing radio supernova monitoring program of nearby starburst galaxies. Very Large Array (VLA) B-configuration snap-shot observations on 2003 Nov. 14 place 3σ upper limits for this source of 0.34 mJy/bm (8.4 GHz) and 1 mJy/bm (15 GHz) at the reported position. MERLIN observations at 5 GHz on 2004 Mar. 9 provide a limit of 0.18 mJy/bm (3σ), some 109 days after its first optical identification. If SN 2004am is a type-II supernova (as suspected from it optical spectrum), a significant delay (up to ∼ 1 year) may be expected before the radio emission peak. Continued radio monitoring of this source is underway with MERLIN and the VLA.

COMET C/2001 Q4 (NEAT) Visual total-magnitude estimates: Apr. 15.40 UT, 4.7 (D. A. J. Sear-

gent, Cowra, N.S.W., naked eye); 18.88, 4.8 (A. Pearce, Nedlands, W. Australia, 8×40 binoculars); 22.89, 4.5 (W. Souza, São Paulo, Brazil, 8×30 binoculars); 23.90, 4.3 (A. Amorim, Florianopolis, Brazil, 7×50 binoculars).

2004 April 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green